# GEORCE: A Fast New Control Algorithm for Computing Geodesics

### v1

@@ A.7 Extension to Finslerian manifolds
The proof of global and local quadratic convergence are completely similar for the Riemannian and Finsler case except that eq. 23 in the Finsler case becomes
@@ B.1 GEORCE for Finsler manifolds
4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
@@ B.2 Sparse Newton Method
$\begin{split}H_{ii}y_{i}&=\left(H_{ii+1}|s_{i}\right),\quad i=1,\dots,T-2,\\ H_{T-1}y_{T-1}&=s_{T-1},\\ \end{split}$ (31)
@@
4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
@@
7:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
@@
Appendix C Manifolds
@@
$\mathbb{S}^{n}$ The n-sphere, $\{x\in\mathbb{R}^{n+1}\;|\;||x||_{2}=1\}$ Stereographic coordinates
$E(n)$ The n-Ellipsoid, $\{x\in\mathbb{R}^{n+1}\;|\;||p\odot x||_{2}=1\}$ Half-axes: $p=\left(0.5,\dots,1\right)$ equally distributed points. Stereographic coordinates
@@
$\mathbb{H}^{2}$ 2 dimensional Hyperbolic Space embedded into Minkowski pace $\mathbb{R}^{3}$ Parameterized by $\left(\alpha,\beta\right)$ with $x=\cosh\alpha,y=\sinh\alpha\cos\beta,z=\sinh\alpha\sin\beta$.
$P(n)$ Paraboloid of dimension $n$ Parameterized by standard coordinates by $\left(x^{1},\dots,x^{n},\sum_{i=1}^{n}x_{i}^{2}\right)$
$\mathcal{P}(n)$ Symmetric positive definite matrices of size $n^{2}$ Embedded into $\mathbb{R}^{n^{2}}$ by the mapping $f(x)=l(x)l(x)^{T}$, where $l:\mathbb{R}^{n(n+1)/2}\rightarrow\mathbb{R}^{n\times n}$ maps $x$ into a lower triangle matrix consisting of the elements in $x$.
@@
where $g_{ij}$ denotes the elements of the Riemannian background metric, $f^{j}$ denotes the $j$th element of the force field, while $v_{0}$ denotes the initial velocity Riemannian length of an object moving with constant velocity $v$ on the surface. For the manifold to be a proper Finsler manifold it is required that $||f||_{g}<v_{0}$.
Appendix D Variational autoencoders
D.1 Architecture
We train a Variational-Autoencoer (VAE) [52], where we use the normal distribution as a variational family to approximate the data distribution. We train a VAE for each of the following datasets
@@ D.1 Architecture
The architectures for the VAE for MNIST data, SVHN data and CelebA data are shown in Table 5, Table 6 and Table 7, respectively. The architecture is similar to [12] with some modifications. All models are trained for $50,000$ epochs on a GPU described in Appendix E.2.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix E Experiments
E.1 Hyper-parameters and methods
@@
$\mathbb{S}^{n}$ $n$ equally spaced points between 00 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
$E(n)$ $n$ equally spaced points between 00 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
@@
E.2 Hardware
All figures have been computed on a HP computer with Intel Core i9-11950H 2.6 GHz 8C, 15.6” FHD, 720P CAM, 32 GB (2$\times$16GB) DDR4 3200 So-Dimm, Nvidia Quadro TI2004GB Descrete Graphics, 1TB PCle NVMe SSD, backlit Keyboard, 150W PSU, 8cell, W11Home 64 Advanced, 4YR Onsite NBD.
@@ E.2 Hardware
The runtime experiments on a CPU have been run on a CPU model XeonE5_2660v3 with $4$ nodes for at most $24$ hours with a maximum memory of $10$ GB. .

### v2

@@ A.2 Extension to Finslerian manifolds
The proof of global and local quadratic convergence is completely similar for the Riemannian and Finsler case except that Eq. 14 in the Finsler case becomes
@@ B.1 GEORCE for Finsler manifolds
4:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
@@ B.2 Sparse Newton Method
[…]
@@
4:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
@@
7:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
@@
Appendix C Choice of initialization curve
Figure 12: The first row shows the estimated curves using GEORCE for the egg-tray, i.e. $f(x,y)=\left(x,y,2\cos x\cos y\right)$, while the second row shows the estimated curves for the sphere for different initializations of the initial curve.
In this section, we show the effect of the choice of initialization curve in GEORCE. In Algorithm 1 and the experiments, the initial curve is a straight line between the start and end point. However, the initial curve is not restricted to this and can in principle be any curve assuming that it is defined in the local chart. To illustrate the effect of the initial curve, we will consider the egg-tray with the parametrization $f(x,y)=\left(x,y,2\cos x\cos y\right)$, and the sphere for two antipodal points on the equator. For the egg-tray we consider setting the initial curve randomly as
$x_{t}^{(0)}=t\frac{b-a}{T}+a+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}\left(0,I\right),$
while we consider for the sphere the random initial curve.
$x_{t}^{(0)}=t\frac{b-a}{T}+a+\boldsymbol{1}\mathrm{Sign}\left(\epsilon\right),\quad\epsilon\sim\mathcal{N}\left(0,1\right),$
For each random initial curve, we estimate the energy-minimizing curve using GEORCE and show the result in Fig. 12 for $100$ random initializations. The first row shows the results for the egg-tray, where the rightmost figure is the distribution of length for the computed curves. We display the corresponding curves in the same color as in the histogram. In general, we see that for a small set of curves GEORCE has found a local minimum, where the curves move on the inside of the egg-tray rather than on the outside of the egg-tray, which appears to be the global minimum. This illustrates how GEORCE is guaranteed to converge to a local minimum and not a global minimum, and that this can be checked by expecting different initialization curves. For the sphere, there are infinitely many length-minimizing curves, and since we add the sign of a standard normal distributed variable for the random curves, we see that GEORCE tends to move in the direction of the closest local minimum.
Appendix D Manifolds
@@
[…]
@@
$\mathbb{H}^{2}$ 2 dimensional Hyperbolic Space embedded into Minkowski pace $\mathbb{R}^{3}$ - Parameterized by $\left(\alpha,\beta\right)$ with $x=\cosh\alpha,y=\sinh\alpha\cos\beta,z=\sinh\alpha\sin\beta$.
$P(n)$ Paraboloid of dimension $n$ - Parameterized by standard coordinates by $\left(x^{1},\dots,x^{n},\sum_{i=1}^{n}x_{i}^{2}\right)$
$\mathcal{P}(n)$ Symmetric positive definite matrices of size $n^{2}$ - Embedded into $\mathbb{R}^{n^{2}}$ by the mapping $f(x)=l(x)l(x)^{T}$, where $l:\mathbb{R}^{n(n+1)/2}\rightarrow\mathbb{R}^{n\times n}$ maps $x$ into a lower triangle matrix consisting of the elements in $x$.
@@
where $g_{ij}$ denotes the elements of the Riemannian background metric, $f^{j}$ denotes the $j$th element of the force field, while $v_{0}$ denotes the initial velocity Riemannian length of an object moving with constant velocity $v$ on the surface. For the manifold to be a proper Finsler manifold it is required that $f||_{g}<v_{0}$.
Appendix E Variational autoencoders
E.1 Architecture
We train a Variational-Autoencoder (VAE) [52], where we use the normal distribution as a variational family to approximate the data distribution. We train a VAE for each of the following datasets
@@ E.1 Architecture
The architectures for the VAE for MNIST data, SVHN data and CelebA data are shown in Table 10, Table 11 and Table 12, respectively. The architecture is similar to [12] with some modifications. All models are trained for $50,000$ epochs on a GPU described in Appendix F.2.
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=1,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=16,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=1,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=3,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
@@
$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=3,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=4,\mathrm{bias}=\mathrm{False}\right)$
@@ Appendix F Experiments
F.1 Hyper-parameters and methods
@@
$\mathbb{S}^{n}$ $n$ equally spaced points between 0 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
$E(n)$ $n$ equally spaced points between 0 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
@@
F.2 Hardware
[…]
@@ F.2 Hardware
The runtime experiments on a CPU have been run on a CPU model XeonE5_2660v3 with $1$ nodes for at most $24$ hours with a maximum memory of $10$ GB.